A Cell phone company offers two plans to its subscribers. At the time new subscribers sign up they are asked to provide some demographic information. The mean yearly income for a sample of 42 subscribers to plan A is $57,000 with a standard Deviation of $6,000. For a sample of 50 subscribers to plan B the mean income is $63,000 with a standard deviation of $4,900. At a significance of 0.05 is there a difference in between the two plans?
Ho
H1
Alpha:
CV:
What Formula?

Answer by stanbon(75887)

A Cell phone company offers two plans to its subscribers. At the time new subscribers sign up they are asked to provide some demographic information.
The mean yearly income for a sample of 42 subscribers to plan A is $57,000 with a standard Deviation of $6,000.
For a sample of 50 subscribers to plan B the mean income is $63,000 with a standard deviation of $4,900.
At a significance of 0.05 is there a difference in between the two plans?
-------------
Comment: I don't think you are comparing plans; you are comparing income
of plan purchasers.
------------------------------
Ho: u(A)=u(B)
H1: u(A) is not equal to u(B)
----------------------
Alpha:5%
--------------
CV:+-1.96
--------------
What Formula?
Test statistic for 2-Sample T-test:
t(57000-63000) = (-6000)/sqrt[(6000^2/42)+(4900^2/50)]=-5.188..
------------
Conclusion: since test stat is in the rejection interval, Reject Ho.
Income of plan A and plan B purchasers are not the same.
=================
Cheers,
Stan H.